OpenStudy (anonymous):

Simplify the expression (4x – 3)(2x2 – 7x + 1)

OpenStudy (littlenugget):

>:D i luv these problems!!!!! hehe okie dokie here we go =)

OpenStudy (anonymous):

haha ok littlenugget

OpenStudy (littlenugget):

You are first going to take the first number in the first set of parentheses, which is "4x" and mutliply it which each number in the 2nd set of parentheses (2x2-7x+1) 

OpenStudy (anonymous):

so multiply it all by 4

OpenStudy (littlenugget):

so... 1st one: \[4x \times 2x ^{2}= 8x ^{3}\] 2nd one: \[4x \times (-7x) = -28x\] 3rd one: \[4x \times 1= 4x\] yep you multiply all by "4x" and then you add them together: \([8x ^{3}-28x^{2}+4x -3)\] Note: the -3 was from the 1st parenthese.... so basically you replaced the "4x" with all the numbers you multiplied

OpenStudy (littlenugget):

So now your equation is \[(8x ^{3}-28x^{2}+4x -3)(2x2-7x+1)\]

OpenStudy (littlenugget):

Now remember what you did to the 4x and how you multiplied it by everythin in the 2nd set of parentheses? You are going to do the same with the "-3" can you do that for me? :)

OpenStudy (anonymous):

well I would like to but I am not all that smart

OpenStudy (littlenugget):

You can do it ;D sorry my internet is going all weird on me -_- So anyways: -3 x 2x2 = -6x2 -3 x -7x = 21x (2 negatives = a positive) -3 x 1= -3 so you are replacing the 3 with -6x2+21x-3 So now let's plug it in with the other numbers :) So now here's your equation :D (8x3-28x2+4x-3-6x2+21x-3) but..... we're not done yet! Now we need to simplify!!!!!

OpenStudy (littlenugget):

So now we are going to add "like terms" things that have the same like of exponent on the "x". for example x^3 with x^3; and x^2 with x^2.. So we are going to put all the "like terms" together to make it easier ( i always do this so i make less silly mistakes): \[8x ^{3}-28x^{2}-6x^{2}+4x+21x-3-3\] Now we can simplify :) And the answer is....... \[8x ^{3}-34x ^{2}+25x-6\]

OpenStudy (littlenugget):

Yeah!!! :D

OpenStudy (littlenugget):

If something I said wasn't clear, just let me know and I'll explain it ;)